# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

### v3

@@ Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench Verified, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce two new diagnostic tasks: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art (SoTA) models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. Similar patterns are also observed for the function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks (up to 35% consecutive 5-gram overlap ratio on SWE-Bench Verified and Full, but only up to 18% for tasks in other benchmarks). These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
1 Introduction
@@
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly (Srivastava et al. 2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks (Zhuo et al. 2024; Pan et al. 2024; Bowman and Dahl 2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues (Jimenez et al. 2024; OpenAI 2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress (Anthropic 2025).
@@
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA (Touvron et al. 2023), PaLM (Chowdhery et al. 2022), and Codex (Chen et al. 2021), creating potential overlap between training and evaluation data.
@@
To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs, and patch generation, the ability to generate the corrected version of the buggy code. Both sub-tasks are fundamental to successful issue resolution. For bug localization (Hossain et al. 2024), models should require a genuine understanding of both the problem description and codebase structure to identify relevant files. For patch generation, models need to understand the semantics of the code and, based on the issue description to reason and generate the correct solution. Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability. To systematically investigate these different memorization patterns, we design three diagnostic tasks: file-path identification, function reproduction, and prefix completion. The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. The function reproduction task measures models’ ability to reproduce fixed function code without providing the necessary information for such reproduction, such as the respective function specification. The prefix completion task checks models’ ability to verbatim generate the solutions of tasks given the tasks’ prefixes.
@@
However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns during training, and there lack of a standard baseline for their understanding of general coding abilities. To address this challenge, we propose a cross-benchmark performance analysis approach and use comparative analysis across similar systems as proxies. In our case, we evaluate models across multiple similar benchmarks and repositories, using performance difference across similar tasks as indicators of memorization. Genuine coding skill should lead to similar performance across comparable tasks, while memorization would result in unusually high scores on tasks that likely overlap with the model’s training data.
@@
Our results reveal concerning patterns: on the file-path identification task, SoTA models like o3 (OpenAI 2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
@@ 2 Approach
2.1 File Path Identification Task
@@
2.2 Function Reproduction Task
@@
2.3 Prefix Completion Task
@@ 3.1 Benchmarks
SWE-Bench-C#
SWE-Bench-C# is an internal benchmark that follows SWE-Bench’s construction pipeline and comprises 75 tasks from eleven C# repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description. This benchmark allows us to test whether performance patterns observed on Python-based tasks extend to tasks from different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities. Details are available in the supplementary materials.
RefactorBench
RefactorBench (Gautam et al. 2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from 9 open-source Python repos. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor. For our experiment, we use a subset of 39 tasks from the full dataset. Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repos and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task or repo-specific optimization rather than general coding proficiency.
@@
3.2 Additional experiment settings
@@
Full-SWE-Bench
We randomly sample 200 instances from the full SWE-Bench (Jimenez et al. 2024) dataset, not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these other instances, it may indicate that models have been optimized specifically for the curated subset or have seen similar patterns during training, i.e., instance-specific memorization.
SWE-Bench Extra
@@
Outside-Repo Tasks
@@ 3.3 Metrics
The accuracy is computed as the percentage of predictions that exactly match the GT patch out of the total number of samples in that category. An exact match is counted when the model-generated file path is identical to at least one of the reference GT file path(s).
@@
To quantify the degree of verbatim memorization in model outputs, we compute 5-gram overlap ratio, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the GT code. This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching that respects the occurrence count of each n-gram in the ground truth to prevent inflation from repeated phrases. This metric captures the extent to which models regurgitate exact token sequences. When models show high overlap on some benchmarks but not others, this differential pattern can indicate memorization of specific training data rather than consistent coding ability.
@@
3.4 Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor (OpenAI 2025a). From Anthropic, our suite comprised the Claude family: 3.5 Sonnet (June 20, 2024), 3.7 Sonnet (February 24, 2025), 4.0 Sonnet, and 4.0 Opus, both officially released on May 22, 2025.
@@ 3.4 Models
4 Results
4.1 File Path Identification Accuracy
@@
Instance-Specific memorization
@@
Repository-Bias memorization
@@
Model-Specific Contamination Patterns
@@
Implications for Benchmark Validity
@@
4.2 File Path Identification Accuracy on Filtered Instances
@@
4.3 Function Reproduction Result
@@
4.4 Direct Memorization on SWE-Bench Verified
@@
5 Related Work
5.1 Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval (Jimenez et al. 2024; Zhuo et al. 2024; Xia, Deng, and Zhang 2024). Among them, SWE-Bench has emerged as the most popular evaluation method for LLM’s coding abilities (Jimenez et al. 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al. 2024; Wang et al. 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
5.2 Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on eval data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding (Zhou et al. 2023). To combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmarks. For example, EvoEval (Xia, Deng, and Zhang 2024) uses LLMs to create semantic-altering transformations of HumanEval (Chen et al. 2021) problems, while other methods (Chen et al. 2025) use AST-level mutations to measure how performance changes with structural code changes. Similarly, TaskEval (Tambon et al. 2025) measures prompt sensitivity by creating variations of a task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze model outputs without changing the task itself. These techniques (Li 2023; Ramos et al. 2025) include using perplexity or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if model has likely seen a specific bug or solution before. However, these measures cannot be applied to commercial models as we do not have access to the model’s hidden states. N-gram similarity, while applicable, is a noisy indicator for complex code patches; any functionally correct solution will inevitably share text overlap with ground truth, making it hard to distinguish genuine problem-solving from verbatim memorization.
6 Conclusion
This work presents evidence that current evaluations of LLM coding abilities may be compromised by benchmark contamination and memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific, evidenced by graduated performance decay within SWE-Bench ecosystem (in both file-path identification and function reproduction tasks), and repository-bias memorization, demonstrated by substantial performance drops on external repos (up to 47 percentage points for file-path identification) despite their popularity. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns of SWE-Bench Verified’s data.
Appendix A SWE-Bench C# Details
We show the repository distribution of SWE-Bench C# in Figure 6. This benchmark is useful for tuning agents problem-solving approach for bugfixing, feature addition, and refactoring in the C# and .NET ecosystem.
Appendix B Prompt Templates and Examples
B.1 File-Path Identification Prompt Template
We present the prompt template for the File-Path Identification task in Figure 7. The prompt begins with the repository name, followed by the issue description, which serves as the problem statement for SWE-Bench Verified tasks, and concludes with instructions for the model to reason through the problem and output a single predicted buggy file path.
B.2 SWE-Bench Verified Task Example
B.3 SWE-Bench C# Example
Figure 9 shows the problem statement for instance: ardalis__cleanarchitecture-530 in SWE-Bench C#.
B.4 Function Reproduction Task Prompt Template
Appendix C Additional Experiment Results
C.1 Comparison of 5-Gram Overlap for Buggy V.S. Ground Truth Code Snippet
We also report the difference of 5 gram overlap
where is the 5 gram overlap of and . Positive values indicate the model output more closely matches the patched implementation; negative values indicate closer similarity to the buggy code.
@@ Appendix C Additional Experiment Results
All models, except Claude 3.5 Sonnet and GPT 4.1, produce negative on SWE‑Bench Extra. Because this split is mined largely from post‑cutoff commits, the patches are unlikely to be in training data. The systematic shift toward the buggy reference therefore supports our contamination‑mitigation claim: when the patched code is unseen, models do not memorize it from context.
@@
Taken together, the function reproduction experiment provides a direct, contrastive probe of code memorization: when historical (patched) code is in a model’s training mix, we observe measurable positive ; when the code post‑dates the model (SWE‑Bench Extra), the signal disappears or reverses. Additionally, the high tendency towards the fixed version of code on SWE-Bench Verified and SWE-Bench C# indicates higher possibility of benchmark memorization on these two settings. These results underscore the necessity of freshness‑controlled benchmarks when interpreting coding model performance and memorization risk.

### v4

@@ Abstract.
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench Verified, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability from memorized patterns and other learned artifacts. In this work, we introduce two new diagnostic tasks: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art (SoTA) models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. Similar patterns are also observed for the function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks (up to 35% consecutive 5-gram overlap ratio on SWE-Bench Verified and Full, but only up to 18% for tasks in other benchmarks). These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
1. Introduction
@@
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly (Srivastava et al., 2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks (Zhuo et al., 2024; Pan et al., 2024; Bowman and Dahl, 2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues (Jimenez et al., 2024; OpenAI, 2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress (Anthropic, 2025).
@@
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA (Touvron et al., 2023), PaLM (Chowdhery et al., 2022), and Codex (Chen et al., 2021), creating potential overlap between training and evaluation data.
@@
To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs, and patch generation, the ability to generate the corrected version of the buggy code. Both sub-tasks are fundamental to successful issue resolution. For bug localization (Hossain et al., 2024), models should require a genuine understanding of both the problem description and codebase structure to identify relevant files. For patch generation, models need to understand the semantics of the code and, based on the issue description to reason and generate the correct solution. Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability. To systematically investigate these different memorization patterns, we design three diagnostic tasks: file-path identification, function reproduction, and prefix completion. The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. The function reproduction task measures models’ ability to reproduce fixed function code without providing the necessary information for such reproduction, such as the respective function specification. The prefix completion task checks models’ ability to verbatim generate the solutions of tasks given the tasks’ prefixes.
@@
However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns during training, and there is a lack of a standard baseline for their understanding of general coding abilities. To address this challenge, we propose a cross-benchmark performance analysis approach and use comparative analysis across similar systems as proxies. In our case, we evaluate models across multiple similar benchmarks and repositories, using performance differences across similar tasks as indicators of memorization. Genuine coding skill should lead to similar performance across comparable tasks, while memorization would result in unusually high scores on tasks that likely overlap with the model’s training data.
@@
Our results reveal concerning patterns: on the file-path identification task, SoTA models like o3 (OpenAI, 2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
@@ 2. Approach
2.1. File Path Identification Task
@@
2.2. Function Reproduction Task
@@
2.3. Prefix Completion Task
@@ 3.1.1. SWE-Bench-Verified
3.1.2. SWE-Bench-C#
SWE-Bench-C# is an internal benchmark that follows SWE-Bench’s construction pipeline and comprises 75 tasks from eleven C# repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description. This benchmark allows us to test whether performance patterns observed on Python-based tasks extend to tasks from different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities. We show the repository distribution of SWE-Bench C# in Figure 4. This benchmark is useful for tuning agents problem-solving approach for bugfixing, feature addition, and refactoring in the C# and .NET ecosystem.
Figure 5 shows an example task in the SWE-Bench C# dataset, instance: ardalis__cleanarchitecture-530.
3.1.3. RefactorBench
RefactorBench (Gautam et al., 2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from 9 open-source Python repos. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor. For our experiment, we use a subset of 39 tasks from the full dataset. Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repos and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task or repo-specific optimization rather than general coding proficiency.
@@ 3.1.3. RefactorBench
3.2. Additional experiment settings
@@
3.2.1. Full-SWE-Bench
We randomly sample 200 instances from the full SWE-Bench (Jimenez et al., 2024) dataset, not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these other instances, it may indicate that models have been optimized specifically for the curated subset or have seen similar patterns during training, i.e., instance-specific memorization.
3.2.2. SWE-Bench Extra
@@
3.2.3. Outside-Repo Tasks
@@ 3.3. Metrics
The accuracy (1) is computed as the percentage of predictions that exactly match the GT patch out of the total number of samples in that category. An exact match is counted when the model-generated file path is identical to at least one of the reference GT file path(s).
@@
To quantify the degree of verbatim memorization in model outputs, we compute 5-gram overlap ratio (2), which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the GT code. This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching that respects the occurrence count of each n-gram in the ground truth to prevent inflation from repeated phrases. This metric captures the extent to which models regurgitate exact token sequences. When models show high overlap on some benchmarks but not others, this differential pattern can indicate memorization of specific training data rather than consistent coding ability.
@@
3.4. Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor (OpenAI, 2025a). From Anthropic, our suite comprised the Claude family: 3.5 Sonnet (June 20, 2024), 3.7 Sonnet (February 24, 2025), 4.0 Sonnet, and 4.0 Opus, both officially released on May 22, 2025.
@@ 3.4. Models
3.5. Prompt Templates
3.5.1. File-Path Identification Prompt Template
We present the prompt template for the File-Path Identification task in Figure 6. The prompt begins with the repository name, followed by the issue description, which serves as the problem statement for SWE-Bench Verified tasks, and concludes with instructions for the model to reason through the problem and output a single predicted buggy file path. Here the ‘{0}’ is the repository name, and the ‘{1}’ is the issue description from the task.
3.5.2. Function Reproduction Task Prompt Template
4. Results
4.1. File Path Identification Accuracy
@@
4.1.1. Instance-Specific memorization
@@
4.1.2. Repository-Bias memorization
@@
4.1.3. Model-Specific Contamination Patterns
@@
4.1.4. Implications for Benchmark Validity
@@
4.2. File Path Identification Accuracy on Filtered Instances
@@
4.3. Function Reproduction Result
@@
4.4. Direct Memorization on SWE-Bench Verified
@@
4.5. Comparison of 5-Gram Overlap for Buggy V.S. Ground Truth Code Snippet
We also report the difference of 5 gram overlap (3), where is the 5 gram overlap of and . Positive values indicate the model output more closely matches the patched implementation; negative values indicate closer similarity to the buggy code.
@@
All models, except Claude 3.5 Sonnet and GPT 4.1, produce negative on SWE‑Bench Extra. Because this split is mined largely from post‑cutoff commits, the patches are unlikely to be in training data. The systematic shift toward the buggy reference, therefore, supports our contamination‑mitigation claim: when the patched code is unseen, models do not memorize it from context.
@@
Taken together, the function reproduction experiment provides a direct, contrastive probe of code memorization: when historical (patched) code is in a model’s training mix, we observe measurable positive ; when the code post‑dates the model (SWE‑Bench Extra), the signal disappears or reverses. Additionally, the high tendency towards the fixed version of code on SWE-Bench Verified and SWE-Bench C# indicates a higher possibility of benchmark memorization on these two settings. These results underscore the necessity of freshness‑controlled benchmarks when interpreting coding model performance and memorization risk.
5. Related Work
5.1. Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval (Jimenez et al., 2024; Zhuo et al., 2024; Xia et al., 2024). Among them, SWE-Bench has emerged as the most popular evaluation method for LLM’s coding abilities (Jimenez et al., 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al., 2024; Wang et al., 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
5.2. Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on eval data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding (Zhou et al., 2023). To combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmarks. For example, EvoEval (Xia et al., 2024) uses LLMs to create semantic-altering transformations of HumanEval (Chen et al., 2021) problems, while other methods (Chen et al., 2025) use AST-level mutations to measure how performance changes with structural code changes. Similarly, TaskEval (Tambon et al., 2025) measures prompt sensitivity by creating variations of a task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze model outputs without changing the task itself. These techniques (Li, 2023; Ramos et al., 2025) include using perplexity or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if model has likely seen a specific bug or solution before. However, these measures cannot be applied to commercial models as we do not have access to the model’s hidden states. N-gram similarity, while applicable, is a noisy indicator for complex code patches; any functionally correct solution will inevitably share text overlap with ground truth, making it hard to distinguish genuine problem-solving from verbatim memorization.
6. Conclusion
This work presents evidence that current evaluations of LLM coding abilities may be compromised by benchmark contamination and memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific, evidenced by graduated performance decay within SWE-Bench ecosystem (in both file-path identification and function reproduction tasks), and repository-bias memorization, demonstrated by substantial performance drops on external repos (up to 47 percentage points for file-path identification) despite their popularity. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns of SWE-Bench Verified’s data during training.